What will my investment of R100 500 be worth if my bank pays me simple interest of 9% per annum and I invest my money for 180 days?
1
Expert's answer
2020-11-30T15:57:49-0500


Interest= principle X rate X time

principle = R100,500

rate = 9%

time = 180 days = 180/365 years

Therefore, Interest "=100,500\times\frac{9}{100}\times\frac{180}{365}" = R4,460.55

Amount = principle + interest = R100,500 + R4,460.55 = R104,960.55

The investment will be worth R104,960.55